# Efficient LiDAR Reflectance Compression via Scanning Serialization

### v1

@@ Abstract
Machine Learning, ICML
@@ 1 Introduction
Point cloud reflectance, derived from the intensity of returned laser pulses in LiDAR, serves as a crucial descriptor of the physical properties of surrounding objects. It is crucial in downstream tasks such as object detection and semantic segmentation (Viswanath et al., 2024; Tatoglu & Pochiraju, 2012). For instance, our experiments show that removing the reflectance causes a dramatic drop in pedestrian (and cyclist) detection accuracy incorporating the widely used PointPillar (Lang et al., 2019) detection model, having AP (average precision) from 51.4 (62.8) to 14.1 (34.3) on the KITTI dataset (see Fig. 1), making it impractical for use. While existing neural compression methods have achieved significant performance gains, they mostly focus on the geometry or RGB color attributes (Wang et al., 2025; You et al., 2024; Zhang et al., 2023; Wang et al., 2023), and the compression of LiDAR reflectance is relatively under-explored with uncompetitive efficiency. Such inferiority of reflectance compression is largely due to the challenges posed by the intrinsic characteristics of LiDAR data.
@@
Extensive experiments on benchmark datasets show that SerLiC achieves remarkably high performance, over 2$\times$ volume reduction against the original reflectance data and up to 22% bit rate reduction compared to the state-of-the-art Unicorn method (Wang et al., 2025), while using only 2% of the parameters and 10% of the GPU memory, as illustrated in Fig. 1. Moreover, SerLiC is hardware-friendly, relying solely on simple networks rather than specific convolution libraries (e.g., Minkowski Engine (Choy et al., 2019)) commonly used in existing methods. Our key contributions are summarized as follows:
@@ 3.1 Problem Definition
The serialization process reorganizes the input point cloud $(\mathbf{X},\mathbf{C})$ into $L$ point sequences as follows:
@@ 3.4 Entropy Coding
Thus, the token $\mathbf{F}^{token}_{i}$ is passed through several Mamba blocks. A standard Mamba layer is shown in Figure 3a, and the processing flow can be summarized as follows:
@@
Figure 3: Basic Mamba block and Attention block. “LN” refers to Layer Norm; $\sigma$ denotes SiLU activation; $\otimes$ means Hadamard product; $\oplus$ represents element-wise addition.
Dual Parallelization. SerLiC proposes a dual parallelization strategy at both the sequence and window levels to accelerate the coding process. First, as shown in Fig. 2, SerLiC processes point sequences in parallel, disregarding correlations across sequences. This is justified, as point sequences are scanned by separate laser sensor rotations, inherently lacking strong inter-sequence dependencies. As a result, sequence parallelization does not compromise coding efficiency but greatly enhances processing speed.
@@ 4.1 Datasets
  • nuScenes is a large-scale dataset collected for autonomous driving using Velodyne HDL-32E. It has ten subsets, each containing 85 scenes. For training, we extract the first 100 frames from the first 12 scenes in the first five subsets, resulting in 6,000 frames. For testing, we select the first 90 frames from the first scene of each of the last five subsets, yielding 450 frames numbered as sequences #01 to #05. Also, we quantify them to 1mm geometric precision with 8-bit reflectance.
@@ 4.2 Experimental Details
Training Setting. We implement SerLiC using Python 3.10 and PyTorch 2.5. The model is trained with AdamW optimizer (Loshchilov & Hutter, 2019), using a learning rate of $2\times 10^{-4}$ and a batch size of 64. We employ a cosine annealing strategy (Loshchilov & Hutter, 2017) to gradually reduce the learning rate to $5\times 10^{-5}$. The model is randomly initialized and trained for 25 epochs on testing datasets. For fair comparisons, all experiments are conducted on the same platform, equipped with an NVIDIA RTX 4090 GPU, an Intel Core i9-13900K CPU, and 64GB of memory.
@@ 4.3 Compression Performance
Notice that the compression of L3C2 relies heavily on raw parameters of LiDAR sensor, including the distance from each laser sensor to the LiDAR center and the number of points each line can scan in a full rotation. For datasets lacking such sensor data, such as KITTI and nuScenes, L3C2 cannot function effectively (hence we only provide L3C2 results on Ford). In contrast, the sensor information used in SerLiC is derived directly from the input point cloud, making it operate seamlessly across all datasets. These results are evidence of the superior effectiveness of SerLiC as well as its robustness on various datasets.
@@ 4.5 Ablation Study
Attention-based Coding. Attention represents another prevalent neural structure for sequence modeling, as exemplified by OctAttention (Fu et al., 2022) and Point Transformer v3 (Wu et al., 2024). For a fair comparison, we substitute the original Mamba module with Masked Multi-Head Attention (Vaswani et al., 2017) (detailed in Figure 3b) while maintaining strict consistency in key hyperparameters, such as the number of layers and channels. We observe that using Attention receives similar coding gains to Mamba but higher complexity. Figure 4 illustrates the complexity of Attention-based and Mamba-based SerLiC under different configurations. It is observed that Attention has much higher complexity, in both decoding delay and GPU memory consumption. This is because Attention computes point dependencies within a sequence window $W$ via matrix multiplication, resulting in quadratic complexity $\Theta(W^{2})$. Particularly for decoding which requires point-by-point processing in a window, the complexity increases to $\Theta(W^{3})$.
@@
Figure 4: Comparison of decoding latency and running memory between Attention and Mamba implementations in SerLiC.
@@ 5 Conclusion
This paper presents SerLiC, a serialization-based neural compression framework tailored for LiDAR reflectance attribute. By leveraging scan-order serialization, SerLiC transforms 3D point clouds into 1D sequences, aligning with LiDAR scanning mechanisms and enabling efficient sequential modeling. The Mamba model with physics-informed tokenization further enhances its ability to capture points correlations autoregressively, while maintaining linear-time complexity. Its high efficiency, ultra-low complexity, and strong robustness make it a practical solution for real LiDAR applications. Future work will extend SerLiC to lossy compression for higher compression efficiency while preserving essential information for downstream tasks.
@@
Appendix A LiDAR Working Mechanism
Figure LABEL:fig:supp_lidar illustrates the working mechanism of LiDAR, which helps readers to better understand our method.
Appendix B Testing Dataset Details
Three widely accepted datasets, including KITTI, Ford, and nuScenes, are used as testing datasets in this work. We visualize typical samples from these three datasets in Fig. LABEL:fig:supp_vis_dataset. Accordingly, Table 1 provides detailed information about our testing datasets.
Notably, the testing datasets, KITTI and Ford, are both captured by Velodyne HDL-64E LiDAR and recommended by the international standard committee MPEG. We adhere strictly to the Common Test Conditions (CTC) defined by the MPEG AI-PCC group to quantize samples in these datasets and evaluate compression performance on these two datasets. Since these two datasets are recommended by MPEG, their reliability is well-established, and testing on them effectively reflects the compression capabilities of the evaluated methods.
Furthermore, to verify the generalizability of our work, we additionally evaluate our SerLiC on the nuScenes dataset that uses Velodyne HDL-32E LiDAR.
Overall, these testing datasets used in our work can sufficiently evaluate our performance. Testing on more datasets yields similar results and will not change the conclusion of this work.
@@ Appendix B Testing Dataset Details
Appendix C Complexity and Analysis
@@
C.1 SerLiC Runtime Analysis
The coding process of SerLiC consists of three stages: context construction (CC, executed on CPU), neural network (NN, executed on GPU), and arithmetic coding (AC, executed on CPU). These three stages work following the proposed Dual Parallelization mechanism when processing a frame, including the sequence level and the window level, as illustrated in Fig. LABEL:subfig:supp_dual_parallel. Each sequence is divided into windows of equal size and all windows are processed in parallel.
@@ C.1 SerLiC Runtime Analysis
  • Frame-level Sequential implementation. We implemented a frame-level sequential version of SerLiC, running point cloud frames sequentially on our platform, without any pipeline structure, as illustrated in Fig. LABEL:subfig:supp_wo_pipeline_example. As reported in Table 2 and Fig. LABEL:fig:codec_speed(a)-(b), our SerLiC attains 0.18/0.23 seconds per frame for encoding/decoding, i.e., $>$5/4 fps. SerLiC (light) achieves a faster speed of $>$10 fps (frames per second). Even using the sequential version, our SerLiC achieves state-of-the-art processing speed. Its light version can meet typical demands in applications such as conventional autonomous driving, LiDAR-based mapping, and geospatial surveying.
  • Frame-level pipeline implementation. For accelerate processing, we arrange CC, NN, and AC on three pipeline stages for frame-level parallelism, as illustrated in Fig. LABEL:subfig:supp_w_pipeline_example. As presented in Table 2 and Fig. LABEL:fig:codec_speed(c)-(d), using the three-stage pipelining across frames, SerLiC runs approximately at a speed of 0.09/0.13 seconds per frame for encoding/decoding, i.e., 11/7.7 fps. The lightweight version of SerLiC, SerLiC (light), is even faster, achieving over 30 fps, meeting the real-time requirements of many LiDAR-related applications like real-time obstacle detection in autonomous driving, virtual and augmented reality (VR/AR), and high-fidelity 3D reconstruction.
@@
C.2 Attention Performance and Complexity
As described in the main manuscript, Attention can also be used in SerLiC for context modeling. However, compared to Attention, Mamba is more suited to our SerLiC framework due to its inherent ability to capture point dependencies in a sequence.
@@ C.2 Attention Performance and Complexity
Implementation Details. Given the parallel window size $W$, the computational complexity of self-attention mechanism is $\Theta(W^{2})$ due to the matrix-based multiplication computation between $Q$, $K$, and $V$ in self-attention. As we use the autoregressive coding method, a causal mask with the size of $W\times W$ (as indicated on the right of Fig. 3 of our main manuscript) is applied to ensure that each token in the sequence only uses its preceding tokens (its following tokens are unavailable as they are not coded yet) for context modeling. During the encoding process, since all points are known to the encoder, the model can process all tokens in parallel, and the complexity is accordingly $\Theta(W^{2})$. However, during decoding, to generate an output sequence consistent with the encoding, the model must decode each point sequentially in a window $W$. As such, the attention mechanism results in a cubic time complexity $\Theta(W^{3})$ in decoding.
@@ C.3 L3C2 Analysis and Complexity
L3C2 Introduction. MPEG introduced L3C2 (Sébastien & Jonathan, 2021) in 2021, specifically tailored for the compression of LiDAR point clouds. However, its dependence on detailed sensor configurations (e.g., the precise pitch angles of individual lasers) limits its applicability on various datasets. Specifically, L3C2 requires the following inputs (refer to Fig. 5):
@@
For readers’ convenience, we capture the configuration of L3C2 codec in Fig. 5 to demonstrate the parameters required by L3C2.
Since MPEG provides detailed LiDAR configuration parameters only for the Ford dataset, we are unable to test on the KITTI and nuScenes datasets in this work. Moreover, these parameters are only applicable to raw, unprocessed LiDAR point cloud data. However, the point cloud samples in nuScenes and KITTI datasets have undergone preprocessing before release. As a result, even if their LiDAR parameters could be derived, they would not be compatible with these preprocessed samples.
@@
Appendix D G-PCC Configuration
@@
Appendix E Downstream Tasks
In our introduction, we mention that “…our experiments show that removing the reflectance attribute causes a dramatic drop in pedestrian (and cyclist) AP (average precision) of the widely used PointPillar (Lang et al., 2019) detection model, from 51.4 (62.8) to 14.1 (34.3) on the KITTI dataset (see Fig. 1)…” To support our statement, we offer our full results on the KITTI dataset in Table 6.
It is observed that, without reflectance data, the mean Average Precision (mAP) for the car category decreases from 78.4 to 74.2, while the performance for pedestrian and cyclist categories drops remarkably — from 51.4 (62.8) to 14.1 (34.3). This substantial decline is due to the fact that LiDAR reflectance is highly sensitive to the material properties of pedestrians and cyclists. Consequently, for tasks such as object detection and classification, reflectance data are indispensable for maintaining model’s high accuracy and robustness.
Table 6: Detection results using PointPillar w/ and w/o reflectance on the KITTI dataset
Method Car AP_R40 Pedestrian AP_R40 Cyclist AP_R40 mAP
w/ reflectance 78.396 51.415 62.807 64.206
w/o reflectance 74.213 14.076 34.283 40.857

### v2

@@ Abstract
Point Cloud Compreesion, Lossless Compression, Attribute
@@ 1 Introduction
Point cloud reflectance, derived from the intensity of returned laser pulses in LiDAR, serves as a crucial descriptor of the physical properties of objects. It is widely used in downstream tasks such as object detection and semantic segmentation (Viswanath et al., 2024; Tatoglu & Pochiraju, 2012). For instance, our experiments show that removing the reflectance causes a dramatic drop in pedestrian (and cyclist) detection accuracy incorporating the pre-trained PointPillar (Lang et al., 2019) detection model, having AP (average precision) from 51.4 (62.8) to 14.1 (34.3) on the KITTI dataset (see Figure 1), making it impractical for use. While existing neural compression methods have achieved significant performance gains, they mostly focus on the geometry or RGB color attributes (Wang et al., 2025; You et al., 2024; Zhang et al., 2023; Wang et al., 2023), and the compression of LiDAR reflectance is relatively under-explored with uncompetitive efficiency. This compression gap is from the intrinsic challenges of reflectance data.
@@
Extensive experiments on benchmark datasets show that SerLiC achieves remarkably high performance, over 2$\times$ volume reduction against the original reflectance data and up to 22% bit rate reduction compared to the state-of-the-art Unicorn method (Wang et al., 2025), while using only 2% of the parameters and 10% of the GPU memory, as illustrated in Figure 1. Moreover, SerLiC is hardware-friendly, relying solely on simple networks rather than specific convolution libraries (e.g., Minkowski Engine (Choy et al., 2019)) commonly used in existing methods. Our key contributions are summarized as follows:
@@ 3.1 Problem Definition
The serialization process reorganizes the input point cloud $(\mathbf{X},\mathbf{C})$ into $L$ point sequences as follows (see Figure 2):
@@ 3.4 Entropy Coding
Thus, the token $\mathbf{F}^{token}_{i}$ is passed through several Mamba blocks. A standard Mamba layer is shown in LABEL:subfig:block_mamba, and the processing flow can be summarized as follows:
@@
Dual Parallelization. SerLiC proposes a dual parallelization strategy at both the sequence and window levels to accelerate the coding process. First, as shown in Figure 2, SerLiC processes point sequences in parallel, disregarding correlations across sequences. This is justified, as point sequences are scanned by separate laser sensor rotations, inherently lacking strong inter-sequence dependencies. As a result, sequence parallelization does not compromise coding efficiency but greatly enhances processing speed.
@@ 4.1 Datasets
  • nuScenes is a large-scale dataset collected for autonomous driving using Velodyne HDL-32E. It has ten subsets, each containing 85 scenes. For training, we extract the first 100 frames from the first 12 scenes in the first five subsets, resulting in 6,000 frames. For testing, we select the first 90 frames from the first scene of each of the last five subsets, yielding 450 frames numbered as sequences #01 to #05. Also, we quantize them to 1mm geometric precision with 8-bit reflectance.
@@ 4.2 Experimental Details
Training Setting. We implement SerLiC using Python 3.10 and PyTorch 2.5. The model is trained with AdamW optimizer (Loshchilov & Hutter, 2019), using a learning rate of $2\times 10^{-4}$ and a batch size of 64. We employ a cosine annealing strategy (Loshchilov & Hutter, 2017) to gradually reduce the learning rate to $5\times 10^{-5}$. The model is randomly initialized and trained for 25 epochs for each dataset. For fair comparisons, all experiments are conducted on the same platform, equipped with an NVIDIA RTX 4090 GPU, an Intel Core i9-13900K CPU, and 64GB of memory.
@@ 4.3 Compression Performance
Notice that the compression of L3C2 relies heavily on detailed parameters of LiDAR sensor, including the elevation angle of each laser sensor and their respective distances to the LiDAR center. For datasets lacking such sensor data, such as KITTI and nuScenes, L3C2 cannot function effectively (hence we only provide L3C2 results on Ford). In contrast, SerLiC only requires basic LiDAR information such as the angular resolutions of the elevation and azimuth angles. These results are evidence of the superior effectiveness of SerLiC as well as its robustness on various datasets.
@@ 4.5 Ablation Study
Attention-based Coding. Attention represents another prevalent neural structure for sequence modeling, as exemplified by OctAttention (Fu et al., 2022) and Point Transformer v3 (Wu et al., 2024). For a fair comparison, we substitute the original Mamba module with Masked Multi-Head Attention (Vaswani et al., 2017) (detailed in LABEL:subfig:block_attn) while maintaining strict consistency in key hyperparameters, such as the number of layers and channels. We observe that using Attention receives similar coding gains to Mamba but higher complexity. LABEL:fig:ablation_attention illustrates the complexity of Attention-based and Mamba-based SerLiC under different configurations. It is observed that Attention has much higher complexity, in both decoding delay and GPU memory consumption. This is because Attention computes point dependencies within a sequence window $W$ via matrix multiplication, resulting in quadratic complexity $\Theta(W^{2})$. Particularly for decoding which requires point-by-point processing in a window, the complexity increases to $\Theta(W^{3})$.
@@
Table 7: Quantitative compression gains against G-PCC (RAHT) on the non-rotational dataset
Class Seq. Bpp CR Gain
RAHT SerLiC vs. RAHT
InnovizQC 02 3.71 2.28 -38.54%
03 3.82 2.25 -41.10%
Avg. 3.77 2.27 -39.71%
Table 8: Quantitative compression gains against G-PCC (RAHT) on various density scenarios in KITTI
Class Seq. Bpp CR Gain
RAHT SerLiC vs. RAHT
KITTI City 4.67 3.24 -30.62%
Residential 4.96 3.81 -23.19%
Road 5.41 4.39 -18.85%
Campus 4.86 3.43 -29.42%
Avg. 4.98 3.72 -25.28%
Non-Rotational Adaptation. In addition to the KITTI, Ford, and nuScenes datasets that are captured by rotational LiDAR scanning, we also evaluated SerLiC on the non-rotational InnovizQC dataset provided by MPEG. It provides three sequences, each having 300 frames with 16-bit coordinates and 8-bit reflectance. We used the first sequence for training and the remaining two for testing. Results shown in Table 7 demonstrate SerLiC’s strong generalization capability to non-rotational LiDAR systems.
Scenario Robustness. To further validate the robustness of SerLiC, we provide its detailed performance on KITTI across various scenarios, as reported in Table 8. Results indicate that SerLiC consistently outperforms G-PCC in all scenes. Relatively, SerLiC achieves better performance in City and Campus scenes compared to Residential and Road scenes. This occurs due to dense roadside vegetation interfering with sensor-recorded reflectance characteristics and thus increasing compression difficulty.
@@ 5 Conclusion
This paper presents SerLiC, a serialization-based neural compression framework tailored for LiDAR reflectance attribute. By leveraging scan-order serialization, SerLiC transforms 3D point clouds into 1D sequences, aligning with LiDAR scanning mechanisms and enabling efficient sequential modeling. The Mamba model with physics-informed tokenization further enhances its ability to capture points correlations autoregressively, while maintaining linear-time complexity. Its high efficiency, ultra-low complexity, and strong robustness make it a practical solution for real LiDAR applications. Future work will extend SerLiC to lossy compression for higher compression efficiency.
@@
Acknowledgments
This research was supported by National Natural Science Foundation of China (62171174) and Natural Science Foundation of Jiangsu Province (BK20243038). We are grateful to the anonymous reviewers for comments on early drafts of this paper. We also extend our sincere thanks to the authors of the relevant works used in our comparative studies for providing the latest results for evaluation.
Appendix A Testing Dataset Details
Three widely accepted datasets, including KITTI, Ford, and nuScenes, are used as testing datasets in this work. We visualize typical samples from these three datasets in LABEL:fig:supp_vis_dataset. Accordingly, Table 1 provides their detailed information.
@@ Appendix A Testing Dataset Details
Appendix B Complexity and Analysis
@@
B.1 SerLiC Runtime Analysis
The coding process of SerLiC consists of three stages: context construction (CC, executed on CPU), neural network (NN, executed on GPU), and arithmetic coding (AC, executed on CPU). These three stages work following the proposed Dual Parallelization mechanism when processing a frame, including the sequence level and the window level, as illustrated in LABEL:subfig:supp_dual_parallel. Each sequence is divided into windows of equal size and all windows are processed in parallel.
@@ B.1 SerLiC Runtime Analysis
  • Frame-level sequential implementation. We implemented a frame-level sequential version of SerLiC, running point cloud frames sequentially on our platform, without any pipeline structure, as illustrated in LABEL:subfig:supp_wo_pipeline_example. As reported in Table 2 and LABEL:fig:codec_speed(a)-(b), our SerLiC attains 0.18/0.23 seconds per frame for encoding/decoding, i.e., $>$5/4 fps (frames per second). SerLiC (light) achieves a faster speed of $>$10 fps. Even using the sequential version, our SerLiC achieves state-of-the-art processing speed. Its light version can meet typical demands in applications such as conventional autonomous driving, LiDAR-based mapping, and geospatial surveying.
  • Frame-level pipeline implementation. To accelerate processing, we arrange CC, NN, and AC on three pipeline stages for frame-level parallelism, as illustrated in LABEL:subfig:supp_w_pipeline_example. As presented in Table 2 and LABEL:fig:codec_speed(c)-(d), using the three-stage pipelining across frames, SerLiC runs approximately at a speed of 0.09/0.13 seconds per frame for encoding/decoding, i.e., 11/7.7 fps. The lightweight version of SerLiC, SerLiC (light), is even faster, achieving over 30 fps, meeting the real-time requirements of many LiDAR-related applications like real-time obstacle detection in autonomous driving, virtual and augmented reality (VR/AR), and high-fidelity 3D reconstruction.
@@
B.2 Attention Performance and Complexity
As described in the main manuscript, Attention can also be used in SerLiC for context modeling. However, compared to Attention, Mamba is more suited to our framework due to its inherent ability to capture point dependencies in a sequence.
@@ B.2 Attention Performance and Complexity
Implementation Details. Given the parallel window size $W$, the computational complexity of self-attention mechanism is $\Theta(W^{2})$ due to the matrix-based multiplication computation between $Q$, $K$, and $V$ in self-attention. As we use the autoregressive coding method, a causal mask with the size of $W\times W$ (as indicated in Figure 3b of our main manuscript) is applied to ensure that each token in the sequence only uses its preceding tokens (its following tokens are unavailable as they are not coded yet) for context modeling. During the encoding process, since all points are known to the encoder, the model can process all tokens in parallel, and the complexity is accordingly $\Theta(W^{2})$. However, during decoding, to generate an output sequence consistent with the encoding, the model must decode each point sequentially in a window $W$. As such, the attention mechanism results in a cubic time complexity $\Theta(W^{3})$ in decoding.
@@ B.3 L3C2 Analysis and Complexity
L3C2 Introduction. MPEG introduced L3C2 (Sébastien & Jonathan, 2021) in 2021, specifically tailored for the compression of LiDAR point clouds. However, its dependence on detailed sensor configurations (e.g., the precise pitch angles of individual lasers) limits its applicability on various datasets. Specifically, L3C2 requires the following inputs (refer to Figure 6):
@@
For readers’ convenience, we capture the configuration of L3C2 codec in Figure 6 to demonstrate the parameters required by L3C2. Since MPEG provides detailed LiDAR configuration parameters only for the Ford dataset, we are unable to test on KITTI and nuScenes in this work. Moreover, these parameters are only applicable to raw, unprocessed LiDAR point cloud data. However, the point cloud samples in nuScenes and KITTI datasets have undergone preprocessing before release. As a result, even if their LiDAR parameters could be derived, they would not be compatible with these preprocessed samples.
@@
Table 6: Detection results using the classical detection models w/ and w/o reflectance on the KITTI dataset
Methods Car Ped. Cyc. mAP
Easy Mod. Hard Easy Mod. Hard Easy Mod. Hard Mod.
PointPillar (w/ R) 87.75 78.40 75.18 57.30 51.41 46.87 81.57 62.81 58.83 64.21
PointPillar (w/o R) 83.85 74.21 70.36 21.56 14.08 12.83 47.79 34.28 32.13 40.86
SECOND (w/ R) 90.55 81.61 78.61 55.95 51.15 46.17 82.97 66.74 62.78 66.50
SECOND (w/o R) 87.87 78.92 75.58 40.46 35.66 31.88 70.42 50.64 47.70 55.07
PointRCNN (w/ R) 91.47 80.54 78.05 62.96 55.04 48.56 89.17 70.89 65.64 68.82
PointRCNN (w/o R) 88.03 77.00 72.85 42.90 35.28 30.68 56.58 42.51 40.07 51.60
Appendix C G-PCC Configuration
@@
Appendix D Downstream Tasks
In our introduction, we mention that “…our experiments show that removing the reflectance causes a dramatic drop in pedestrian (and cyclist) detection accuracy incorporating the pre-trained PointPillar (Lang et al., 2019) detection model, having AP (average precision) from 51.4 (62.8) to 14.1 (34.3) on the KITTI dataset (see Figure 1)…” To support our statement, we offer our full results on the KITTI dataset in Table 6.
The results in Table 6 provide a detailed comparison of the detection performance of three classical models—PointPillar, SECOND, and PointRCNN—with and without reflectance data. Note that our experiments are conducted on pre-trained models. The table shows the Average Precision (AP) for three categories: Car, Pedestrian (Ped.), and Cyclist (Cyc.), across different difficulty levels (Easy, Moderate, Hard), as well as the mean Average Precision (mAP) for the Moderate category.
It is observed that, reflectance data is essential for maintaining high accuracy and robustness in object detection tasks, particularly for pedestrians and cyclists. The substantial drop in AP for these categories underscores the importance of preserving reflectance information in point cloud data for downstream applications.